OpenStudy (anonymous):

Solve on the interval (0, 2pi): 1-cos theta =2sqrt3/2 How would I solve this? Thank you :)

OpenStudy (anonymous):

first solve for \(\cos(\theta)\)

OpenStudy (anonymous):

The 1-cos theta is what's confusing me. I can't see how the answer relates to the unit circle

OpenStudy (anonymous):

first of all i am confused by \(\frac{2\sqrt3}{2}\) is that really what it says?

OpenStudy (anonymous):

Sure is :/

OpenStudy (anonymous):

ok \[\frac{2\sqrt3}{2}=\sqrt3\] right ?

OpenStudy (anonymous):

then you get \[1-\cos(\theta)=\sqrt3\] so \[-\cos(\theta)=\sqrt3-1\] and therefore \[\cos(\theta)=1-\sqrt3\]

OpenStudy (anonymous):

after that you are on your own the only way i could solve this is to say \[\theta=\cos^{-1}(1-\sqrt3)\] and use a calculator

OpenStudy (anonymous):

Oh! Okay, thanks for getting me started. I think I can get it from there.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i still think \[\frac{2\sqrt3}{2}\] is a strange thing to write